# A high-order recombination algorithm for weak approximation of stochastic differential equations

### v1

@@ 3.1. Refined error estimation
Then, using the Taylor expansion, there exists a polynomial $q_{U_{k}^{(i)}}$ of order $r+1$ such that
$\sup_{y\in U_{k}^{(i)}}\left|P^{X}_{t}f(y)-q_{U_{k}^{(i)}}(y)\right|\\ \leq\operatorname{rad}\left(U_{k}^{(i)}\right)^{r+1}\sup_{y\in U_{k}^{(i)}}\max_{q_{1}+\dots+q_{N}=r+1}\left|\frac{\partial^{q_{1}}}{\partial x_{1}^{q_{1}}}\cdots\frac{\partial^{q_{N}}}{\partial x_{N}^{q_{N}}}f(y)\right|$ (33)
for $y_{0}\in U_{k}^{(i)}$. By using the property of reduced measure, we have
$\begin{split}&\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\left|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right|\\ &\leq\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\sum_{j=1}^{L_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\left(\max_{k_{1}+\dots+k_{N}=m+1}\left.\nu^{(i)}\right|_{U_{j}^{(i)}}\left(\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f\right)\right.\\ &\qquad\qquad\qquad\left.-\max_{k_{1}+\dots+k_{N}=m+1}\left.\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\right|_{U_{j}^{(i)}}\left(\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f\right)\right)\\ &\leq\sum_{i=1}^{n-1}\sum_{j=1}^{L_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\operatorname{wt}\left(U_{k}^{(i)}\right)\max_{k_{1}+\dots+k_{N}=m+1}\sup_{y\in U_{k}^{(i)}}\left|\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f(y)\right|\end{split}$
Using (32), we have
$\begin{split}&\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\left|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right|\\ &\qquad\leq C_{2}(T)\sum_{i=1}^{n-1}\frac{\sum_{j=1}^{L_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\operatorname{wt}(U_{k}^{(i)})}{(T-t_{i})^{m/2}}\|\nabla f\|_{\infty}\end{split}$
thus. ∎
Example 2 (Refined patch conditon).
For patches $\left\{U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=1,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ that satisfy
@@
In this subsection,we present a recursive algorithm for constructing a set of patches $\left\{U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=1,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ that satisfies (34) from the probability measure $\nu^{(i)}=\sum_{j=1}^{m_{i}}\hat{w}_{j}^{(i)}\delta_{p^{(i)}_{j}}$. To this end, given a measure $\bar{\nu}=\sum_{j=1}^{m}w_{j}\delta_{p_{j}}$ on $\mathbb{R}^{N}$, we define a penalty function $\Psi_{i}$ as
@@
where $U^{\bar{\nu}}=\operatorname{supp}(\bar{\nu})=\left\{p_{j}\right\}_{j=1}^{m}$. The penalty function corresponds to the difference between the radius and the patch condition given in Example 2.
@@
Figures 2 and 3 show an example of patch division. This example is in the case of pricing an Asian option under the Heston model, which we discuss in subsection 4.1, and the number of partitions $n=6$ and the fifth times step $t_{5}=(5/6)T$. Here, the sizes represent the weights of the nodes, the colors represent which patch the node is included in. We can see that nodes with smaller weights, that is smaller size nodes, are divided to one large patch.
Remark 4 (Advantages of the above-described patch division algorithm).
The above algorithm enables a patch division that satisfies the condition in (34), while minimizing the total number of patches as much as possible. Moreover, due to its recursive nature, the patch division can be executed efficiently. This reduces the overall computational cost of the recombination method.
@@ 4.1. Convergence of Discretisation error
$\begin{split}&dY^{(1)}(t,x)=\mu Y^{(1)}(t,x)\;\mathrm{d}t+Y^{(1)}(t,x)\sqrt{Y^{(2)}(t,x)}\;\mathrm{d}B^{(1)}(t),\\ &dY^{(2)}(t,x)=\alpha(\theta-Y^{(2)}(t,x))\;\mathrm{d}t+\beta\sqrt{Y^{(2)}(t,x)}\left(\rho\;\mathrm{d}B^{(1)}(t)+\sqrt{1-\rho^{2}}\;\mathrm{d}B^{(2)}(t)\right)\end{split}$
@@
which is weaker than (40). This condition for the positivity can be derived the discretisation formula presented in [24]. The payoff of the Asian call option on this asset with maturity $T$ and strike $K$ is $\max(Y^{(3)}(T,x)/T-K,0)$ where
@@
The Heston model do not satisfy the UH condition which we assume in Theorem 5. Nevertheless, the numerical result demonstrates that our algorithm works for the Heston model case. In fact, the following considerations generally justify the numerical result. In the Heston model, the positivity of the discretised process by the N–V method ensure that $\operatorname{supp}(\nu^{(i)})\subset\mathbb{R}^{3}_{>0}=\left\{(y_{1},y_{2},y_{3})\;|\;y_{1},y_{2},y_{3}>0\right\}$, where $\nu^{(i)}$ is an intermediate measure defined in (11). This indicates that for any patch division $\left\{U_{k}^{(i)}\right\}_{k=1,\dots,l_{i}}$ of $\operatorname{supp}(\nu^{(i)})$ there exists an $\varepsilon>0$ such that $\bigcup_{k=1,\dots,l_{i}}U_{k}^{(i)}\subset\left\{(y_{1},y_{2},y_{3})\;|\;y_{1},y_{2},y_{3}>\varepsilon\right\}$. Moreover, in the Heston model, $\left\{W_{1},W_{2},[W_{0},W_{1}]\right\}$ spans >$\mathbb{R}^{3}_{>0}$, thus, Theorem 2 of [12] implies that for each patch $U_{k}^{(i)}$ there exists a $C_{U_{k}^{(i)}}$ such that (32) holds, and
@@
4.2. Impact of patch division algorithm on recombination error
To investigate the effect of patch division on recombination error, we conduct further validation. We introduce a scaling parameter $\kappa$ to the penalty function as follows.
$\Psi^{(\kappa)}(U)=\text{rad}(U)-\kappa(T-t_{i})^{{m}/({2(m+1)})}s_{i}^{{1}/{2}}\left(\operatorname{wt}(U)\right)^{{1}/(m+1)}.$ (42)
For a fixed the number of partition $n$, by varying $\lambda$, we calculate the recombination errors
$\left|E_{\mu_{x,\Delta,\mathcal{U},\kappa}^{(n)}}\left[f\left(Y^{(n)}(T,x)\right)\right]-E_{\lambda_{x,\Delta,\mathcal{U}}^{(n)}}\left[f\left(Y(T,x)\right)\right]\right|,$
namely the difference between the expectations under the recombined measure $\mu_{x,\Delta,\mathcal{U},\kappa}^{(n)}$ constructed by using the penalty function (42) and the measure without recombining $\lambda_{x,\Delta,\mathcal{U}}^{(n)}$, that is the whole tree calculation defined by (9), for different number of support. We compare our patch division algorithm (Recursive_Refcond) with the concentric circle dividing (CC), the random dividing (Rand), and the recursive path dividing with Litterer–Lyons condition (Recursive_LLcond). Note that the theoretical patch division condition is not necessarily satisfied in CC and Rand.
Figure 6 shows the recombination errors in pricing the Asian option in the Heston model, as the same setting as in the previous subsection. Here, we fix the number of partitions $n=8$. It seems that convergence efficiencies of each patch division are as follows:
${\rm Recursive\_Refcond}>{\rm Recursive\_LLcond}>>{\rm Rand}>{\rm CC}.$
Figure 6. Total Error (Heston)
@@ 4.2. Impact of patch division algorithm on recombination error
where $h$ is the height, $r$ is the radius, and $(x_{c},y_{c})$ is the centre of cone. We set the height as $h=\frac{10}{r^{2}}$so that the volume of the cones remain constant. In this validation, We consider the following 80 cone payoffs in this coordinate.
@@
Let $\nu$ be an induced probability measure on a patch $U$, defined in (14). $\mathcal{N}_{\nu}$ denotes the number of nodes included in the patch $U$. Let $G=\{g_{1},g_{2},\dots,g_{\mathcal{M}_{G}}\}$ be a set of test functions, where $\mathcal{M}_{G}$ is the number of test functions and each $g_{i}:\tilde{\Omega}\rightarrow\mathbb{R}$ is a measurable function. Hereinafter, $x^{(j)}$ denotes the $j$th coordinate of $x$, that is $x=\left(x^{(1)},\dots,x^{(N)}\right)$.
@@
      $\begin{split}&d_{i,l+1}=\frac{\left(v_{(i-1),(l+1)}\right)^{(e_{i})}}{\left(v_{(i-1),1}\right)^{(e_{i})}},\\ &v_{i,l}=v_{(i-1),(l+1)}-d_{i,l+1}v_{(i-1),1}\in\mathbb{R}^{\mathcal{N}_{\nu}},\\ &\left(v_{i,l}\right)^{(j)}=\left\{\begin{array}[]{c}\left(v_{i,l}\right)^{(j)}\;\quad\text{for $j=1,\dots,e_{i}-1$}\\ \left(v_{i,l}\right)^{(j+1)}\;\quad\text{for $j=e_{i},\dots,\mathcal{N}_{\nu}-i$}\end{array}\right.\end{split}$ (47)
@@
Let $0=t_{0}<t_{1}<\dots<t_{n}=T$ be a partition of time $[0,T]$.
@@
Output : $E_{\mu_{t_{n}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}}f(X(T,x_{0}))$
@@ 3 Struct Node contains
        /* Realized value $\hat{X}(t_{i},x_{0}))(\omega_{j}\in\Omega_{i}$) */
@@
        /* Weight (a value of measure at this node $P\left(\omega_{j}\right)$) */
@@ 21       $value\leftarrow\text{Payoff}(nextleavesleaves,f)$;
1 Function OneStepFwd($leaves,i,n,T,V_{0},\dots,V_{d}$):
@@ 2       $nextleaves\leftarrow\text{array of {Node}}$;
3       forall $node$ in $leaves$ do
              /* Generate RVs $\{\eta_{i}\}$ and solve ODEs */
@@
5                   nextval $\leftarrow$ Solve ODEs along with $V_{0},\dots,V_{d}$;
                    /* Using (23) */

### v2

@@ 3.1. Refined error estimation
Then by the Taylor expansion there exists a polynomial $q_{U_{k}^{(i)}}$ of order $r+1$ such that
$\sup_{y\in U_{k}^{(i)}}\left|P^{X}_{t}f(y)-q_{U_{k}^{(i)}}(y)\right|\\ \leq\operatorname{rad}\left(U_{k}^{(i)}\right)^{r+1}\max_{q_{1}+\dots+q_{N}=r+1}\left|\frac{\partial^{q_{1}}}{\partial x_{1}^{q_{1}}}\cdots\frac{\partial^{q_{N}}}{\partial x_{N}^{q_{N}}}f(y_{0})\right|$ (36)
for $y_{0}\in U_{k}^{(i)}$. From the property of reduced measure, it holds that
$\begin{split}&\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\left|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right|\\ &\leq\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\sum_{k=1}^{l_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\left(\max_{k_{1}+\dots+k_{N}=m+1}\left.\nu^{(i)}\right|_{U_{k}^{(i)}}\left(\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f\right)\right.\\ &\qquad\qquad\qquad\left.-\max_{k_{1}+\dots+k_{N}=m+1}\left.\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\right|_{U_{k}^{(i)}}\left(\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f\right)\right)\\ &\leq\sum_{i=1}^{n-1}\sum_{k=1}^{l_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\operatorname{wt}\left(U_{k}^{(i)}\right)\max_{k_{1}+\dots+k_{N}=m+1}\sup_{y\in U_{k}^{(i)}}\left|\frac{\partial^{k_{1}}}{\partial x_{1}^{k_{1}}}\cdots\frac{\partial^{k_{N}}}{\partial x_{N}^{k_{N}}}P^{X}_{T-t_{i}}f(y)\right|\end{split}$
Hence from the inequality(35)
$\begin{split}&\sum_{i=1}^{n-1}\sup_{x_{0}\in\mathbb{R}^{N}}\left|\nu^{(i)}\left(P^{X}_{T-t_{i}}f\right)-\mu_{t_{i}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}\left(P^{X}_{T-t_{i}}f\right)\right|\\ &\qquad\leq C_{2}(T)\sum_{i=1}^{n-1}\frac{\sum_{k=1}^{l_{i}}\operatorname{rad}\left(U_{k}^{(i)}\right)^{m+1}\operatorname{wt}(U_{k}^{(i)})}{(T-t_{i})^{m/2}}\|\nabla f\|_{\infty}\end{split}$
holds thus. ∎∎
Example 2 below stands in the same relation to Theorem 5 as Example 1 does to Theorem 4. It thus provides a refined version of the patch condition given in Example 1. This condition is hereafter referred to as the weighted LL patch condition, sometimes abbreviated as the WLL patch condition.
Example 2 (WLL patch conditon).
For a set of patch divisions $\left\{U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=1,\dots,\sharp\Delta,\\ k=1,\dots,l_{i}\end{subarray}}$ that satisfy
@@ 3.2. Patch division algorithm
In this subsection, we present a recursive algorithm for constructing a set of patche divisions $\left\{U_{k}^{(i)}\right\}_{\begin{subarray}{c}i=1,\dots,\sharp\Delta\\ k=1,\dots,l_{i}\end{subarray}}$ that satisfies (37), based on the probability measure $\nu^{(i)}=\sum_{j=1}^{m_{i}}\hat{w}_{j}^{(i)}\delta_{p^{(i)}_{j}}$. To this end, we consider a measure $\bar{\nu}=\sum_{j=1}^{m}w_{j}\delta_{p_{j}}$ on $\mathbb{R}^{N}$ and define a penalty function $\Psi_{i}$ as
@@
where $U^{\bar{\nu}}=\operatorname{supp}(\bar{\nu})=\left\{p_{j}\right\}_{j=1}^{m}$. The penalty function corresponds to the difference between the radius and the WLL patch condition.
@@
Implementation advantages.
The above algorithm produces a patch division that satisfies the condition in (37), while minimising the total number of patches as far as possible. Moreover, due to its recursive nature, the division can be performed efficiently. This contributes to reducing the overall computational cost of the recombination method.
Figures 3 and 3 show an example of patch division. This example corresponds to the pricing of an Asian option under the Heston model, discussed in subsection 4.1, with the number of partitions $n=6$ and the fifth time step $t_{5}=(5/6)T$. Here the node size indicates its weight, and the colour shows the patch to which it belongs. We observe that nodes with smaller weights—that is, those represented by smaller-sized circles—are grouped into a single large patch.
@@ 4.1. Convergence of Discretisation error
$\begin{split}&dY^{(1)}(t,x)=\mu Y^{(1)}(t,x)\;\mathrm{d}t+Y^{(1)}(t,x)\sqrt{Y^{(2)}(t,x)}\;\mathrm{d}B^{(1)}(t),\\ &dY^{(2)}(t,x)=\alpha(\theta-Y^{(2)}(t,x))\;\mathrm{d}t\\ &\hskip 108.12047pt+\beta\sqrt{Y^{(2)}(t,x)}\left(\rho\;\mathrm{d}B^{(1)}(t)+\sqrt{1-\rho^{2}}\;\mathrm{d}B^{(2)}(t)\right)\end{split}$
@@
which is weaker as a condition on $\beta$ than (43). This condition for the positivity can be derived the discretisation formula presented in [24]. The payoff of the Asian call option on this asset with maturity $T$ and strike $K$ is $\max(Y^{(3)}(T,x)/T-K,0)$ where
@@
The Heston model do not satisfy the UH condition which we assume in Theorem 5. Nevertheless, the numerical result demonstrates that our algorithm works for the Heston model case. In fact, the following considerations generally justify the numerical result. In the Heston model the positivity of the discretised process by the N–V method ensures that $\operatorname{supp}(\nu^{(i)})\subset\left(\mathbb{R}_{>0}\right)^{3}\coloneqq\left\{(y_{1},y_{2},y_{3})\;|\;y_{1},y_{2},y_{3}>0\right\}$, where $\nu^{(i)}$ is an intermediate measure defined in (12). This indicates that for any patch division $\left\{U_{k}^{(i)}\right\}_{k=1,\dots,l_{i}}$ of $\operatorname{supp}(\nu^{(i)})$ there exists an $\varepsilon>0$ such that $\bigcup_{k=1,\dots,l_{i}}U_{k}^{(i)}\subset\left\{(y_{1},y_{2},y_{3})\;|\;y_{1},y_{2},y_{3}>\varepsilon\right\}$. Moreover, in the Heston model, the set of vectors $\left\{W_{1}(x),W_{2}(x),[W_{0},W_{1}](x)\right\}$ spans the linear space $\mathbb{R}^{3}$ for each $x\in\left(\mathbb{R}_{>0}\right)^{3}$. Thus, restricting the domain of $P_{t}^{X}f$ to each patch $U_{k}^{(i)}$ ensures the UH condition, which implies the UFG condition. Hence, Theorem 2 of [12] yields a constant $C_{U_{k}^{(i)}}$ such that (35) holds, and
@@
4.2. Effect of the patch division algorithm on recombination error
To investigate how patch division affects recombination error, we conduct further validation. The comparison includes the specific algorithm described in the previous subsection, which we hereafter refer to as Recursive_WLL, as well as other alternative patch division strategies.
We introduce a scaling parameter $\lambda>0$ into the penalty function defined as follows:
$\Psi^{(\lambda)}_{i}(U)=\text{rad}(U)-\lambda(T-t_{i})^{{m}/({2(m+1)})}{s_{i}}^{{1}/{2}}\left(\operatorname{wt}(U)\right)^{{1}/(m+1)}.$ (45)
For a fixed number of partitions $n$, we vary $\lambda$ and compute the recombination errors
$\left\lvert\mu_{T}^{\left({\rm NV},\Delta,{\rm Rec},\lambda\right)}\left(f\right)-\mu_{T}^{\left({\rm NV},\Delta\right)}\left(f\right)\right\rvert,$ (46)
that is the difference between the expectation under the recombined measure $\mu_{T}^{\left({\rm NV},\Delta,{\rm Rec},\lambda\right)}$ constructed using the penalty function (45), and that under the ‘un-recombined’ measure $\mu_{T}^{\left({\rm NV},\Delta\right)}$, defined as the full tree-based measure defined by (10) without applying recombination.
We compare the algorithm Recursive_WLL with concentric circle dividing (CC), random dividing (Rand), and recursive path dividing with the Litterer–Lyons condition (Recursive_LL). Note that the theoretical patch division condition is not necessarily satisfied in CC and Rand.
Figure 6 shows the recombination errors in pricing the Asian option under the Heston model, using the same setting as in the previous subsection. Here we fix the number of partitions at $n=8$. Among the methods compared, Recursive_WLL achieves the smallest recombination error, followed by Recursive_LL. The random dividing method yields moderate accuracy, while CC results in the largest error among the four.
Figure 6. Recombination Error (Heston)
@@ 4.2. Effect of the patch division algorithm on recombination error
where $h$ is the height, $r$ is the radius, and $(x_{c},y_{c})$ is the centre of cone. We set the height as $h={10}r^{-2}$ so that the volume of the cones remain constant. In this validation, We consider the following 80 cone payoffs in this coordinate.
@@
Let $\nu$ be an induced probability measure on a patch $U$, defined in (15). $\mathcal{N}_{\nu}$ denotes the number of nodes included in the patch $U$. Let $G=\{g_{1},g_{2},\dots,g_{\mathcal{M}_{G}}\}$ be a set of test functions, where $\mathcal{M}_{G}$ is the number of test functions and each $g_{i}:\tilde{\Omega}\rightarrow\mathbb{R}$ is a measurable function. Hereafter, $x^{(j)}$ denotes the $j$th coordinate of $x$, that is $x=\left(x^{(1)},\dots,x^{(N)}\right)$.
@@
      $\begin{split}d_{i,l+1}&=\frac{\left(v_{(i-1),(l+1)}\right)^{(e_{i})}}{\left(v_{(i-1),1}\right)^{(e_{i})}},\\ v_{i,l}&=v_{(i-1),(l+1)}-d_{i,l+1}v_{(i-1),1}\in\mathbb{R}^{\mathcal{N}_{\nu}},\\ \left(v_{i,l}\right)^{(j)}&=\left\{\begin{array}[]{c}\left(v_{i,l}\right)^{(j)}\;\quad\text{for $j=1,\dots,e_{i}-1$}\\ \left(v_{i,l}\right)^{(j+1)}\;\quad\text{for $j=e_{i},\dots,\mathcal{N}_{\nu}-i$}\end{array}\right.\end{split}$ (51)
@@
Let $\Delta=\{0=t_{0}<\dots<t_{n}=T\}$ be a partition of time $[0,T]$.
@@
Output : $E_{\mu_{t_{n}}^{\left({\rm Alg},\Delta,{\rm Rec}\right)}}f(X^{({\rm NV},\Delta,\rm{Rec})}(T))$
@@ 3 Struct Node contains
        /* Realized value $X^{({\rm NV},\Delta,\rm{Rec})}(t_{i})(\omega_{j}\in\Omega_{i}$) */
@@
        /* Weight (Prob. measure of this node $P\left(\omega_{j}\right)$/
@@ 21       $value\leftarrow\text{Payoff}(nextleavesleaves,f)$;
Function OneStepFwd($leaves,i,n,T,V_{0},\dots,V_{d}$);
@@ 2       $nextleaves\leftarrow\text{array of {Node}}$;
       forall $node$ in $leaves$;
        /* Generate RVs $\{\eta_{i}\}$, solve ODEs */
3       do
@@
                   nextval $\leftarrow$ Solve ODEs along with $V_{0},\dots,V_{d}$;
                    /* (24) */